Find the following values, using the equations, and then work the problems using a financial calculator to check your answers. Disregard rounding differences. (Hint: If you are using a financial calculator, you can enter the known values and then press the appropriate key to find the unknown variable. Then, without clearing the TVM register, you can "override" the variable that changes by simply entering a new value for it and then pressing the key for the unknown variable to obtain the second answer. This procedure can be used in parts b and d, and in many other situations, to see how changes in input variables affect the output variable.) Do not round intermediate calculations. Round your answers to the nearest cent.

  1. An initial $400 compounded for 1 year at 4.5%.

    $  

  2. An initial $400 compounded for 2 years at 4.5%.

    $  

  3. The present value of $400 due in 1 year at a discount rate of 4.5%.

    $  

  4. The present value of $400 due in 2 years at a discount rate of 4.5%.

    $  
Answer #1

a. PV =100
Rate =4.5%
Number of Years =1
FV =PV*(1+r)^n =100*(1+4.5%)^1 =104.50
Using Financial Calculator
N=1;I/Y=4.5%;PV =-100;CPT FV =104.50

b. PV =100
Rate =4.5%
Number of Years =2
FV =PV*(1+r)^n =100*(1+4.5%)^2 =109.20
N=2;I/Y=4.5%;PV =-100;CPT FV =109.20

c. FV =400
Number of Years =1
Rate =4.5%
PV=FV/(1+r)^n =400/(1+4.5%)^1=382.78
N=2;I/Y=4.5%;FV =-400;CPT PV =382.78

d. FV =400
Number of Years =2
Rate =4.5%
PV=FV/(1+r)^n =400/(1+4.5%)^2=366.29
N=2;I/Y=4.5%;FV =-400;CPT PV=366.29
Answer #2

a. PV =100
Rate =4.5%
Number of Years =1
FV =PV*(1+r)^n =100*(1+4.5%)^1 =104.50
Using Financial Calculator
N=1;I/Y=4.5%;PV =-100;CPT FV =104.50

b. PV =100
Rate =4.5%
Number of Years =2
FV =PV*(1+r)^n =100*(1+4.5%)^2 =109.20
N=2;I/Y=4.5%;PV =-100;CPT FV =109.20

c. FV =400
Number of Years =1
Rate =4.5%
PV=FV/(1+r)^n =400/(1+4.5%)^1=382.78
N=2;I/Y=4.5%;FV =-400;CPT PV =382.78

d. FV =400
Number of Years =2
Rate =4.5%
PV=FV/(1+r)^n =400/(1+4.5%)^2=366.29
N=2;I/Y=4.5%;FV =-400;CPT PV=366.29